OpenStudy (anonymous):

Choose the correct procedure for calculating (f + g)(−2) if f(x) = 13x2 and g(x) = 7x − 5. Calculate f(−2) and g(−2), then add them together. Calculate f(−2) and g(−2), then write the results as an ordered pair. Calculate f(−2) and g(−2), then subtract them. Calculate g(−2), plug the result into f(x) and simplify.

OpenStudy (anonymous):

first answer

OpenStudy (anonymous):

(f+g)(-2) = f(-2) + g(-2)

OpenStudy (anonymous):

wats answer

OpenStudy (anonymous):

f(-2) = 13(-2)^2 =13*4=52 g(-2) =7(-2) -5= -14-5 =-19 f(-2) + g(-2) = 52-19

OpenStudy (anonymous):

33

OpenStudy (anonymous):

did u read the answer choices abocve

OpenStudy (anonymous):

yes that why I told you the first answer is the corect procedure evaluate and the adding both